40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6278     Entry time: Tue Feb 14 08:22:27 2012
Author: steve 
Type: Update 
Category: SUS 
Subject: sus damping restored 

ITMX, PRM and BS watchdogs are tripped. They were restored.

Stable MC was disabled so I can use MC_REFL 1 W beam to measure green glass .

ELOG V3.1.3-